# NetBench: A Large-Scale and Comprehensive Network Traffic Benchmark Dataset for Foundation Models

### 15 Mar 2024 v1

@@ Abstract
In computer networking, network traffic refers to the amount of data transmitted in the form of packets between internetworked computers or systems. Monitoring and analyzing network traffic is crucial for ensuring the performance, security, and reliability of a network. However, a significant challenge in network traffic analysis is to process diverse data packets including both ciphertext and plaintext. While many methods have been adopted to analyze network traffic, they often rely on different datasets for performance evaluation. This inconsistency results in substantial manual data processing efforts and unfair comparisons. Moreover, some data processing methods may cause data leakage due to improper separation of training and test data. To address these issues, we introduce NetBench, a large-scale and comprehensive benchmark dataset for assessing machine learning models, especially foundation models, in both traffic classification and generation tasks. NetBench  is built upon seven publicly available datasets and encompasses a broad spectrum of 20 tasks, including 15 classification tasks and 5 generation tasks. Furthermore, we evaluate eight State-Of-The-Art (SOTA) classification models and two generative models using our benchmark. The results show that foundation models significantly outperform the traditional deep learning methods in traffic classification. We believe NetBench  will facilitate fair comparisons among various approaches and advance the development of foundation models for network traffic. Our benchmark is available at https://github.com/WM-JayLab/NetBench.
@@ 1 Introduction
Over the past decades, machine learning-based methods have been extensively developed for network traffic analysis. Earlier studies  [7, 6, 19, 16, 8] primarily assessed the proposed methods on a single dataset encompassing at most 2 classification tasks. Despite achieving notable performance, these studies suffer from disparate data processing pipelines and insufficiently comprehensive evaluations. To address this issue, some recent works  [4, 5, 22] have evaluated performance across multiple datasets and a broader range of classification tasks. For instance, ET-BERT [5], a pre-trained foundation model for network traffic, was trained on 7 datasets covering 5 downstream tasks. However, the data processing techniques in existing works are often customized for themselves. Moreover, some studies randomly split the extracted packets from the same flow into training and testing dataset for packet-level evaluation, which may cause data leakage since the packets from same flow are strongly correlated. Additionally, certain research practices [20, 3, 18] only focused on traffic generation for network simulation rather than traffic classification. In summary, there is a lack of evaluation against a unified and comprehensive traffic benchmark that encompasses both network traffic classification and generation, making it hard to fairly compare performance.
@@
To address existing limitations, we propose NetBench, a large-scale and comprehensive network traffic benchmark covering 7 distinct datasets, featuring 15 classification tasks and 5 generation task, as illustrated in Table 1. For the classification benchmark, we construct 15 tasks across 7 datasets, including ISCXVPN 2016 [1], ISCXTor 2016 [2], USTC-TFC 2016 [17], Cross-Platform [15] dataset contains both Android and IOS applications, CIRA-CIC-DoHBrw 2020 [10], and CIC IoT Dataset 2023 [11]. The specific tasks for each dataset are outlined in table 1. For the generation benchmark, we devise five tasks tailored to essential header fields  [20] at the packet level, including IP addresses (Source/Destination), port number (Source/Destination) and packet length over the same 7 datasets above. This benchmark can help create valid network packets for network simulation.
@@
To prevent data leakage, we first split the original network traffic data into training, validation, and testing sets, from which we extract flows and packets. Then, we anonymize sensitive header fields and employ a unified hexadecimal encoding to standardize different data formats. Moreover, our benchmark provides both flow-level and packet-level evaluations for each dataset, accommodating various input types. To the best of our knowledge, NetBench  is the first network traffic benchmark that covers a wide range of tasks through a unified data processing method. Furthermore, we evaluate eight SOTA models on 15 classification tasks and two generative models on 5 generation tasks using our benchmark. We believe that our benchmark has laid a solid foundation for evaluating network traffic models fairly, which will significantly contribute to the development of foundation models for network traffic.
@@
  • Comprehensive Network Traffic Benchmark. we first create a large-scale and comprehensive network traffic benchmark from 7 distinct datasets with 20 tasks. Our benchmark covers diverse data types and a wide range of tasks, which provides fair comparisons of existing machine learning models in both traffic classification and generation.
  • Unified Data Processing. we introduce a uniform hexadecimal encoding method to process diverse network traffic data with anonymization in sensitive header fields. Our data processing method unifies data pre-processing, data standardization, and data segmentation, which can offer a fair evaluation opportunity and save a lot of human labor.
  • Benchmarking SOTA Methods. we compare the performance of eight SOTA classifiers and two generative models using our benchmark. The evaluation results demonstrate the superiority of foundation models over other approaches in traffic classification.
@@ 2 Related Work
Evaluation on Multiple datasets. To achieve a more comprehensive evaluation, recent works have assessed their models on more datasets covering a variety of tasks. For instance, TSCRNN [4] was evaluated on 3 public datasets following the pre-processing pipelines established in [19, 16]. More recently, Transformer-based foundation models like ET-BERT [5] have been evaluated on 5 classification tasks using 6 public datasets and 1 dataset they created. However, the construction of their packet-level evaluation datasets through random sampling raises the potential issue of data leakage, as highly correlated packets from the same flow could appear in both the training and testing datasets. Additionally, YaTC [22] pre-trained on 4 datasets and fine-tuned on 5 public datasets to test its generalization ability. Nevertheless, it faces a similar risk of data leakage for packet-level evaluation due to its use of a data processing methodology akin to ET-BERT. Meanwhile, NetShare [20] evaluated its generative performance across six datasets, although it is not designed for traffic classification tasks.
@@ 3.1 Data Preparation
To create a large-scale and comprehensive benchmark, we first collect raw data samples from 7 publicly available datasets, as detailed in table 1. Then our next step is to convert them into a standardized format as model input. Fig. 2 shows the proposed pipeline of data preparation, which consists of three steps: Data Pre-Processing, Data Standardization, and Data Segmentation.
  • Data Pre-Processing. Network traffic data encompasses flows of network packets containing privacy-sensitive information. As network traffic packets are captured and saved in Packet Capture (PCAP) files, we firstly segments PCAP files into distinct set for training, validation, and testing. This strategy ensures that high-correlated packets from the same flow will not be simultaneously present in both training and testing data. This effectively reduces the risk of data leakage. Subsequently, we extract flows containing multiple packets from these PCAP files according to distinct combinations of IP addresses, port numbers and protocols  [12]. To protect data privacy, we further anonymize each packet by masking the source/destination IP addresses and port numbers (replace their values with 0). This not only preserves data integrity but also protects data privacy.
@@ 4 Experiments
Table 3: Comparison results of traffic classification tasks from tasks 1 to 8 (flow: Flow-level, pkt: Packet-level, AC: Accuracy, F1: F1-score). Pre-trained models like ET-BERT and YaTC outperform other traditional deep learning methods.
@@
Table 4: Comparison results of traffic classification tasks from tasks 9 to 15 (flow: Flow-level, pkt: Packet-level, AC: Accuracy, F1: F1-score). Pre-trained models like ET-BERT and YaTC outperform other traditional deep learning methods.
@@ 4.1 Experimental Settings
SOTA Models. We assess the performance of 8 major SOTA open-source models on  NetBench to ensure a comprehensive and unbiased comparison. Specifically, they include 7 classification models, DataNets [16], FS-Net [7], BiLSTM_ATTN [19], DeepPacket [8], TSCRNN [4], ET-BERT [5], YaTC [22], as well as two competitive generative model, STAN [18] and Netshare [20]. The detailed experimental settings of each model are described below. For FS-Net, we follow its setting that only evaluate its performance in classification tasks at flow-level. For DataNets, DeepPacket, TSCRNN, and BiLSTM_ATTN, we assess their performance on all classification tasks at both flow level and packet level. Regarding pre-training models such as ET-Bert and YaTC, we employ the released pre-trained weights and fine-tune them with the entirety of prepared training dataset. For generative models, we following the same setting in prior work [20] to randomly sample consecutive packets from each dataset, ensuring a consistent evaluation for future studies.
@@ 4.2 Evaluation Results
Classification Tasks. Table 3 and 4 illustrate the comparison results of the 8 SOTA models for traffic classification using NetBench. From the evaluation results, we can conclude with three main insights: (1) Traditional deep learning methods, such as DataNets, DeepPacket, FS-Net, TSCRNN and BiLSTM_ATTN, exhibit limitations in generalizing to new tasks, with a noticeable tendency to bias classifications towards dominant classes. This is evidenced by the f1-score that is consistently below 0.5 with low recall score due to data imbalance. (2) Pre-trained language models like ET-BERT and YaTC significantly outperform these traditional approaches, showcasing their superior prediction accuracy and generalization ability. (3) Although a flow with multiple packets inherently contains richer information compared to a single packet, pre-trained foundation models trained at flow level do not surpass those trained on packet-level in tasks 3, 8, and 10. This is attributed to the constraints on the input length for pre-trained models as the input flows exceeding a specified length will be truncated.
@@
5 Conclusions
We introduced NetBench, a large-scale and comprehensive benchmark for network traffic analysis, which addressed the critical issue of fair evaluation and comparison among different methods. Characterized by its comprehensive design, NetBench  included a total of 20 evaluation tasks across 7 datasets through a unified data processing approach. Furthermore, we evaluated some SOTA models on our benchmark. The experimental results underscored the benefits of pre-training foundation models, which demonstrated superior performance across a wide range of classification tasks. These observations highlighted the significant potential of foundation models to revolutionize network traffic analysis. For traffic generation, very few foundation models have investigated in this field, which remains to explore in the future. In short, we believe that our benchmark will promote advancements in foundation models for network traffic through fair and comprehensive comparisons.

### 19 Mar 2024 v2

@@ Abstract
In computer networking, network traffic refers to the amount of data transmitted in the form of packets between internetworked computers or Cyber-Physical Systems. Monitoring and analyzing network traffic is crucial for ensuring the performance, security, and reliability of a network. However, a significant challenge in network traffic analysis is to process diverse data packets including both ciphertext and plaintext. While many methods have been adopted to analyze network traffic, they often rely on different datasets for performance evaluation. This inconsistency results in substantial manual data processing efforts and unfair comparisons. Moreover, some data processing methods may cause data leakage due to improper separation of training and testing data. To address these issues, we introduce the NetBench, a large-scale and comprehensive benchmark dataset for assessing machine learning models, especially foundation models, in both network traffic classification and generation tasks. NetBench is built upon seven publicly available datasets and encompasses a broad spectrum of 20 tasks, including 15 classification tasks and 5 generation tasks. Furthermore, we evaluate eight State-Of-The-Art (SOTA) classification models (including two foundation models) and two generative models using our benchmark. The results show that foundation models significantly outperform the traditional deep learning methods in traffic classification. We believe NetBench will facilitate fair comparisons among various approaches and advance the development of foundation models for network traffic. Our benchmark is available at https://github.com/WM-JayLab/NetBench.
@@ 1 Introduction
Over the past decades, machine learning-based methods have been extensively developed for network traffic analysis. Earlier studies [6, 5, 19, 15, 7] primarily assessed the proposed methods on a single dataset encompassing at most 2 classification tasks. Despite achieving notable performance, these studies suffer from disparate data processing pipelines and insufficiently comprehensive evaluations. To address this issue, some recent works [3, 4, 22] have evaluated performance across multiple datasets and a broader range of classification tasks. For instance, the foundation model ET-BERT [4] was pre-trained on 7 datasets to learn intrinsic representations and then fine-tuned for 5 downstream tasks. However, the data processing techniques in existing works are often customized for themselves. Moreover, some studies randomly split the extracted packets from the same flow into training and testing dataset for packet-level evaluation, which may cause data leakage since the packets from same flow are strongly correlated. Additionally, certain research practices [20, 18] only focused on traffic generation for network simulation rather than traffic classification. In summary, there is a lack of evaluation against a unified and comprehensive network benchmark that encompasses both network traffic classification and generation, making it hard to fairly compare performance.
@@
To address existing limitations, we propose NetBench, a large-scale and comprehensive network traffic benchmark covering 7 distinct datasets, featuring 15 classification tasks and 5 generation tasks, as illustrated in Table 1. For the classification benchmark, we construct 15 tasks across 7 datasets, including ISCXVPN 2016 [1], ISCXTor 2016 [2], USTC-TFC 2016 [17], Cross-Platform [14] dataset containing both Android and IOS applications, CIRA-CIC-DoHBrw 2020 [9], and CIC IoT Dataset 2023 [10]. The specific tasks for each dataset are outlined in table 1. For the generation benchmark, we devise five tasks tailored to essential header fields [20] at the packet level, including IP addresses (Source/Destination), port number (Source/Destination) and packet length over the same 7 datasets above. This benchmark can help create valid network packets for network simulation.
@@
To prevent data leakage, we first split the original network traffic data into training, validation, and testing sets, from which we extract flows and packets. Then, we anonymize sensitive header fields and employ a unified hexadecimal encoding to standardize different data formats. Moreover, our benchmark provides both flow-level and packet-level evaluations for each dataset, accommodating various input types. To the best of our knowledge, NetBench is the first network traffic benchmark that covers a wide range of tasks through a unified data processing method. Furthermore, we evaluate eight SOTA models (including two foundation models, ET-BERT [4] and YaTC [22]) on 15 classification tasks and two generative models on 5 generation tasks using our benchmark. We believe that our benchmark has laid a solid foundation for evaluating network traffic models fairly, which will significantly contribute to the development of foundation models for network traffic.
@@
  • Comprehensive Network Traffic Benchmark. We first create a large-scale and comprehensive network traffic benchmark from 7 distinct datasets with 20 tasks. Our benchmark offers diverse data types for foundation model training and a wide range of downstream tasks for fair evaluation in both network traffic classification and generation.
  • Unified Data Processing. We introduce a uniform hexadecimal encoding method to process diverse network traffic data with anonymization in sensitive header fields. Our data processing method unifies data pre-processing, data standardization, and data segmentation, which can offer a fair evaluation opportunity and save human labor.
  • Benchmarking SOTA Methods. We compare the performance of eight SOTA classifiers (including two foundation models) and two generative models using our benchmark. The evaluation results demonstrate the superiority of foundation models over other approaches in traffic classification.
@@ 2 Related Work
Evaluation on Multiple datasets. To achieve a more comprehensive evaluation, recent works have assessed their models on more datasets covering a variety of tasks. For instance, TSCRNN [3] was evaluated on 3 public datasets following the pre-processing pipelines established in [19, 15]. More recently, Transformer-based foundation models like ET-BERT [4] pre-trained on 6 public and 1 collected dataset to learn network traffic representations and then fine-tuned for 5 classification tasks. However, the construction of their packet-level evaluation datasets through random sampling raises the potential issue of data leakage, as highly correlated packets from the same flow could appear in both the training and testing datasets. Additionally, the foundation model YaTC [22] pre-trained on 4 datasets for better representation and fine-tuned on 5 public datasets to test its generalization ability. Nevertheless, it faces a similar risk of data leakage for packet-level evaluation due to its use of a data processing methodology akin to ET-BERT. Meanwhile, NetShare [20] evaluated its generative performance across six datasets, although it is not designed for traffic classification tasks.
@@ 3.1 Data Preparation
To create a large-scale and comprehensive benchmark, we first collect raw data samples from 7 publicly available datasets, as detailed in table 1. Then, our next step is to convert them into a standardized format as model input. Fig. 2 shows the proposed pipeline of data preparation, which consists of three steps: Data Pre-Processing, Data Standardization, and Data Segmentation.
  • Data Pre-Processing. Network traffic data encompasses flows of network packets containing privacy-sensitive information. As network traffic packets are captured and saved in Packet Capture (PCAP) files, we firstly segments PCAP files into distinct set for training, validation, and testing. This strategy ensures that high-correlated packets from the same flow will not be simultaneously present in both training and testing data. This effectively reduces the risk of data leakage. Subsequently, we extract flows containing multiple packets from these PCAP files according to distinct combinations of IP addresses, port numbers and protocols [11]. To protect data privacy, we further anonymize each packet by masking the source/destination IP addresses and port numbers (replaced with 0). This not only preserves data integrity but also protects data privacy.
@@ 4 Experiments
Table 3: Comparison results of traffic classification tasks from tasks 1 to 8 (flow: Flow-level, pkt: Packet-level, AC: Accuracy, F1: F1-score). Foundation models like ET-BERT and YaTC outperform other traditional deep learning methods.
@@
Table 4: Comparison results of traffic classification tasks from tasks 9 to 15 (flow: Flow-level, pkt: Packet-level, AC: Accuracy, F1: F1-score). Foundation models like ET-BERT and YaTC outperform other traditional deep learning methods.
@@ 4.1 Experimental Settings
SOTA Models. We assess the performance of 8 major SOTA open-source models on NetBench to ensure a comprehensive and unbiased comparison. Specifically, they include 7 classification models, DataNet [15], FS-Net [6], BiLSTM_ATTN [19], DeepPacket [7], TSCRNN [3], ET-BERT [4], YaTC [22], as well as two competitive generative models, STAN [18] and Netshare [20]. Note that, ET-BERT and YaTC are foundation models which could be pre-trained and then fine-tuned on different classification tasks. The detailed experimental settings of each model are described below. For FS-Net, we follow its setting that only evaluate its performance in classification tasks at flow-level. For DataNet, DeepPacket, TSCRNN, and BiLSTM_ATTN, we assess their performance on all classification tasks at both flow level and packet level. Regarding ET-Bert and YaTC, we employ the released pre-trained weights and fine-tune them with the entirety of prepared training dataset. For generative models, we following the same setting in prior work [20] to randomly sample consecutive packets from each dataset, ensuring a consistent evaluation for future studies.
@@ 4.2 Evaluation Results
Classification Tasks. Table 3 and 4 illustrate the comparison results of the 8 SOTA models for traffic classification using NetBench. From the evaluation results, we can conclude with three main insights: (1) Traditional deep learning methods, such as DataNet, DeepPacket, FS-Net, TSCRNN and BiLSTM_ATTN, exhibit limitations in generalizing to new tasks, with a noticeable tendency to bias classifications towards dominant classes. This is evidenced by the f1-score that is consistently below 0.5 with low recall score due to data imbalance. (2) Foundation models like ET-BERT and YaTC significantly outperform these traditional approaches, showcasing their superior prediction accuracy and generalization ability. (3) Although a flow with multiple packets inherently contains richer information compared to a single packet, foundation models trained at flow level do not surpass those trained on packet-level in tasks 3, 8, and 10. This is attributed to the constraints on the input length for foundation models as the input flows exceeding a specified length will be truncated.
@@
5 Conclusion
We introduced NetBench, a large-scale and comprehensive benchmark for network traffic analysis, which addressed the critical issue of fair evaluation and comparison among different methods. Characterized by its comprehensive design, NetBench included a total of 20 evaluation tasks across 7 datasets through a unified data processing approach. Furthermore, we evaluated some SOTA models on our benchmark. The experimental results underscored the benefits of foundation models, which demonstrated superior performance across a wide range of classification tasks. These observations highlighted the significant potential of foundation models to revolutionize network traffic analysis. For traffic generation, very few foundation models have investigated in this field. This remains to explore in the future as foundation models have the potential to excel in generation tasks as they do in classification [16]. Besides, longer input length is needed for foundation models to better understand the increasingly complex network traffic. In short, future research could leverage NetBench to train foundation models in a convenient and straightforward manner for a variety of downstream network traffic classification and generation tasks. We believe that our benchmark will promote advancements in foundation models for network traffic through fair and comprehensive comparisons.